OpenStudy (anonymous):

Find the domain: \[f(x)=\frac{ 1 }{ \frac{ 4 }{ x-1 } -2}\]

OpenStudy (primeralph):

Well, you don't want 4/(x-1) to equal infinity or 2, so solve for x in those values.

OpenStudy (jdoe0001):

\(\bf f(x)=\cfrac{ 1 }{ \frac{ 4 }{ x-1 } -2} \\ \quad \\ f(x)=\cfrac{ 1 }{ \frac{ 4 }{ {\color{brown}{ +1}}-1 } -2}\to \cfrac{ 1 }{ \frac{ 4 }{ {\color{brown}{ 0}} } -2}\quad {\color{brown}{ undefined}} \\ \quad \\ or \\ \quad \\ f(x)=\cfrac{ 1 }{ {\color{brown}{ +2}} -2}\to \cfrac{ 1 }{ {\color{brown}{ 0}} }\quad {\color{brown}{ undefined}}\)

OpenStudy (anonymous):

thank you both